# A Spatiotemporal, Quasi-experimental Causal Inference Approach to Characterize the Effects of Global Plastic Waste Export and Burning on Air Quality Using Remotely Sensed Data

### v3

@@ 4.1 Statistical Analysis Results
Figure 4 shows the estimated ADT curve and 95% bootstrap confidence interval (CI) from the main analysis. Positive ADT estimates provide evidence of post-China ban increases in PM2.5 concentrations at GPW sites, compared to those expected under business-as-usual. Thus, our results indicate that at open dump sites in Indonesia, the monthly PM2.5 in 2018-2019 was generally higher than expected under business-as-usual (except for dump sites nowhere near port activity), with the effect getting as large as 1.68 $\mu g/m^{3}$ (95% CI = [0.72, 2.48]) depending on the port proximity. As expected, we see smaller effects at locations with lower port proximity – the pointwise confidence intervals include zero until $\delta=0.38$. Interestingly, the effect curve peaks in the middle and decreases for high values of port proximity, including zero again at $\delta=0.81$ – our intuition is that this could reflect smaller increases in dumping/burning in densely populated areas very close to ports, where there is likely more government oversight and less open space. Integrating across all values of the estimated effect curve yields an average increase at dump sites of 0.86 $\mu g/m^{3}$ attributable to China’s ban on plastic imports, corresponding to a 3.4% increase in overall PM2.5 in 2018-2019 compared to 2012-2017. If we only consider dump sites where the confidence intervals do not include zero, the average increase is 1.51 $\mu g/m^{3}$, corresponding to a 5.3% increase at those sites.
While similarity between the effect curve from our proposed multiply-robust method and that produced using an outcome regression alone (Figure 4) suggests a fairly strong outcome model, variations in the tails highlight the importance of guarding against misspecification of the outcome model by incorporating propensity scores. We also see that the generalized propensity score estimation is more uncertain, resulting in substantially wider confidence intervals for the multiply-robust method than for the outcome regression alone. For a more extreme comparison, the confounder-naive results are shown in Figure S6: the mean effect ranges from -6.4 to 8.5 $\mu g/m^{3}$. This divergence from our main result is unsurprising, but demonstrative.
@@ 5 Discussion
In this case study, we observed a 3.4% increase in PM2.5 at dump sites in Indonesia in 2018-2019 compared to 2012-2017. This effect is on a similar scale to the results of Unfried and Wang (2022), who assessed the domestic impacts of China’s plastic waste import ban and estimated an average reduction of 3.7 $\mu g/m^{3}$, corresponding to 8.5% reduction of overall PM2.5 concentrations in China as a result of their 2018 ban. It makes sense that our observed effect would be smaller, given that Indonesia was only one of the countries to which waste that would otherwise have gone to China was diverted. Unfried and Wang (2022) similarly did not observe an anticipation effect in mid-2017, when China announced its upcoming ban to the World Trade Organization.
@@
This study has a few limitations which merit remarks. First, while we are focused on the air quality impact of China’s 2018 ban and resultant increase in plastic waste in Indonesia, the observed difference between 2018-2019 and 2012-2017 might be affected by another policy that was implemented around the same time (2017-2018): Indonesia’s National Action Plan on Marine Plastic Debris. The goal of this policy was to reduce the amount of plastic waste ending up in the ocean – with limited emphasis on reducing the amount of plastic waste being generated (Arifin et al., 2023); by 2021 Indonesia achieved an estimated 15% reduction of marine plastic debris from 2017 levels (Wang and Karasik, 2022). Ethnographic studies from around the world have found that campaigns to raise awareness about plastic pollution often contribute to increased open burning, as communities seek to cut down on litter and deal with waste that has been collected (Pathak et al., 2024). Hence, it is possible that these two policies were interacting to generate the observed increase in PM2.5 at dump sites in Indonesia. Our analysis showing no pre-trends effect when using 2017 as the “treated” year guards against this possibility somewhat, however the effect of the National Action Plan could lag or increase after 2017.
@@ Appendix C Simulations Exploring the Spatial Weighted Bootstrap
In this appendix, our primary aim is to evaluate the characteristics of the confidence intervals generated by our Spatial Weighted Bootstrap procedure, relative to those generated by a Non-spatial Weighted Bootstrap. Here, calculation of the ADT curve’s bias and mean absolute error (MAE) is an intermediate step to better understand the confidence intervals, as Hettinger et al. (2025a) present extensive investigation of a very similar ADT estimator.
@@ C.4 Simulation Results and Discussion
To elucidate the extent to which under-coverage on the tails is driven by the finite sample bias of LLKR, we did as suggested by Hettinger et al. (2025a) and evaluated the coverage after subtracting out the average bias. Figure S18 indicates that especially at higher values of port proximity, the LLKR bias is playing a substantial role in the coverage for both the Spatial and Non-Spatial Weighted Bootstrap approaches. At lower values of port proximity, the coverage gets pulled closer to 95% for all scenarios except that with increased spatial range.

### v4

@@ 4.1 Statistical Analysis Results
Figure 4 shows the estimated ADT curve and 95% bootstrap confidence interval (CI) from the main analysis. Positive ADT estimates provide evidence of post-China ban increases in PM2.5 concentrations at GPW sites, compared to those expected under business-as-usual. Thus, our results indicate that at open dump sites in Indonesia, the monthly PM2.5 in 2018-2019 was generally higher than expected under business-as-usual (except for dump sites nowhere near port activity), with the effect getting as large as 1.68 $\mu g/m^{3}$ (95% CI = [0.72, 2.48]) depending on the port proximity. As expected, we see smaller effects with lower port proximity – the pointwise confidence intervals include zero until $\delta=0.38$. Interestingly, the effect curve peaks in the middle and decreases for high values of port proximity, including zero again at $\delta=0.81$ – our intuition is that this could reflect smaller increases in dumping/burning in densely populated areas very close to ports, where there is likely more government oversight and less open space. Integrating across all values of the estimated effect curve yields an average increase at dump sites of 0.86 $\mu g/m^{3}$ attributable to China’s ban on plastic imports, corresponding to a 3.3% increase in overall PM2.5 in 2018-2019 compared to 2012-2017. For illustration of the heterogeneity: if we only consider values of port proximity for which the confidence intervals do not include zero, the average increase is 1.51 $\mu g/m^{3}$.
@@
Similarity between the effect curve from our proposed multiply-robust method and that produced using an outcome regression alone (Figure 4) suggests a fairly strong outcome model. Differences between these two methods at the tails may indicate that the generalized propensity score is doing its job (guarding against misspecification of the outcome model) or may reflect LLKR having instability at the extremes, as seen in our simulations (Appendix C). For a more extreme comparison, the confounder-naive results are shown in Figure S6: the mean effect ranges from -6.4 to 8.5 $\mu g/m^{3}$. This divergence from our main result is unsurprising, but demonstrative.
@@ 5 Discussion
In this case study, we observed a 3.3% average increase in PM2.5 at dump sites in Indonesia in 2018-2019 compared to 2012-2017, attributable to China’s plastic waste import ban. This effect is on a similar scale to the results of Unfried and Wang (2022), who assessed the domestic impacts of China’s ban and estimated an average reduction of 3.7 $\mu g/m^{3}$, corresponding to 8.5% reduction of overall PM2.5 concentrations in China as a result of their 2018 ban. It makes sense that our observed effect would be smaller, given that Indonesia was only one of the countries to which waste that would otherwise have gone to China was diverted. Unfried and Wang (2022) similarly did not observe an anticipation effect in mid-2017, when China announced its upcoming ban to the World Trade Organization.
@@
This study has a few limitations which merit remarks. First, while we are focused on the air quality impact of China’s 2018 ban and resultant increase in plastic waste in Indonesia, the observed difference between 2018-2019 and 2012-2017 might be affected by another policy that was implemented around the same time (2017-2018): Indonesia’s National Action Plan on Marine Plastic Debris. The goal of this policy was to reduce the amount of plastic waste ending up in the ocean – with limited emphasis on reducing the amount of plastic waste being generated (21); by 2021 Indonesia achieved an estimated 15% reduction of marine plastic debris from 2017 levels (Wang and Karasik, 2022). Ethnographic studies from around the world have found that campaigns to raise awareness about plastic pollution often contribute to increased open burning, as communities seek to cut down on litter and deal with waste that has been collected (Pathak et al., 2024). Hence, it is possible that these two policies were interacting to generate the observed increase in PM2.5 at dump sites in Indonesia. Our analysis showing no pre-trends effect when using 2017 as the “treated” year guards against this possibility somewhat, however the effect of the National Action Plan could lag or increase after 2017.
@@ Appendix C Simulations Exploring the Spatial Weighted Bootstrap
In this appendix, our primary aim is to evaluate the characteristics of the confidence intervals generated by our Spatial Weighted Bootstrap procedure, relative to those generated by a Non-spatial Weighted Bootstrap. Here, calculation of the ADT curve’s bias and mean absolute error (MAE) is an intermediate step to better understand the confidence intervals, as Hettinger et al. (2025b) present extensive investigation of a very similar ADT estimator.
@@ C.4 Simulation Results and Discussion
To elucidate the extent to which under-coverage on the tails is driven by the finite sample bias of LLKR, we did as suggested by Hettinger et al. (2025b) and evaluated the coverage after subtracting out the average bias. Figure S18 indicates that especially at higher values of port proximity, the LLKR bias is playing a substantial role in the coverage for both the Spatial and Non-Spatial Weighted Bootstrap approaches. At lower values of port proximity, the coverage gets pulled closer to 95% for all scenarios except that with increased spatial range.